All Values of  ln(z) . Find all values and graph some of them in the complex plane. (a) Inſe) (b) (6)In(-7) (c) In (4+3i)

Thanks in advance.
Answer #1

bacz). let 2=2+2y tano a= reosco, y = rfno s nyert 2 = rose firme o -kan (S/M). 2= reasotino z = reio In 12) = ln crero) - I= ln (7) +12011 In 1-7)= ln (7) +i ent , n=0,1,2,3,-- put nao = ln (-7)= ln (7) +20) en(-a)= ln (7) in the principal value of